# Study of $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$

### v1

@@ Abstract
Using a sample of $(10087\pm 44)\times 10^{6}$ $J/\psi$ events accumulated with the BESIII detector, we analyze the decays $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ ($l=e$ or $\mu$) via the process $J/\psi\rightarrow\gamma\eta$. The branching fraction of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ is measured to be $\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})=(3.07\pm 0.12_{\rm{stat.}}\pm 0.19_{\rm{syst.}})\times 10^{-4}$. No signal events are observed for the $\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ decay, leading to an upper limit on the branching fraction of $\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-})<4.0\times 10^{-7}$ at the 90% confidence level. Furthermore, the $CP$-violation asymmetry parameter is found to be $\mathcal{A}_{CP}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})=(-4.04\pm 4.69_{\rm{stat.}}\pm 0.14_{\rm{syst.}})\%$, showing no evidence of $CP$-violation with current statistics. Additionally, we extract the transition form factor from the decay amplitude of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$. Finally, axion-like particles are searched for via the decay $\eta\rightarrow\pi^{+}\pi^{-}a,a\rightarrow e^{+}e^{-}$, and upper limits on this branching fraction relative to that of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ are presented as a function of the axion-like particle mass in the range $5-200\ \mathrm{MeV}/c^{2}$.
@@ I INTRODUCTION
The decays $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ ($l=e$ or $\mu$) involve contributions from the box anomaly of quantum chromodynamics (QCD) and have been studied using various models, such as the hidden gauge Borasoy and Nissler (2007), the chiral unitary approach (Unitary $\chi$PT) Borasoy and Nissler (2007), and the vector meson dominance (VMD) Petri (2010) models. Table 1 displays the theoretical predictions and experimental results for the branching fractions ($\mathcal{B}$) of the $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ decays. Electromagnetic decays of mesons depend on the coupling of photons to the electric charge distribution of the quark fields. This electromagnetic meson structure is described by the transition form factor (TFF), which is a function of the momentum transfer. More importantly, the TFF determines the size of certain hadronic quantum corrections in the anomalous magnetic moment of the muon, $(g-2)_{\mu}$ Danilkin et al. (2019).
@@
CMD-2 Akhmetshin et al. (2001) $3.7^{+2.5}_{-1.8}\pm 3.0$..
WASA Berlowski et al. (2008) $4.3^{+0.2}_{-1.6}\pm 0.4$ $<3.6\times 10^{5}$
KLOE Ambrosino et al. (2009) $2.68\pm 0.09\pm 0.07$..
In the year 2000, the KTeV experiment observed a significant $CP$ asymmetry in the distribution of the T-odd angle $\varphi$ in $K^{0}_{L}\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay Alavi-Harati et al. (2000), where $\varphi$ is the angle between the $e^{+}e^{-}$ decay plane and the $\pi^{+}\pi^{-}$ decay plane in the $K^{0}_{L}$ center-of-mass system. However, so far there is no experimental indication of $CP$-violation in flavor-conserving reactions. The $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay provides an opportunity to search for an analogous $CP$-violation asymmetry in the angle distribution between the $e^{+}e^{-}$ and $\pi^{+}\pi^{-}$ decay planes. This has been explored in $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay in the KLOE experiment, and the $CP$-violation asymmetry is determined to be $\mathcal{A}_{CP}=(-0.6\pm 2.5\pm 1.8)\%$ Ambrosino et al. (2009), which is consistent with zero.
The hadronic decay channels of the $\eta$ meson could also provide signals of the QCD axion, a dark photon or other axion-like particles (ALPs) Landini and Meggiolaro (2020); Alves and Gonzàlez-Solís (2024) that couple with hadrons. Hints of a new light bosonic state with mass around $17\ \mathrm{MeV/c^{2}}$ Krasznahorkay et al. (2016, 2019) were observed by the ATOMKI Collaboration via measuring the angle between $e^{+}e^{-}$ pairs Buttazzo et al. (2021); Ellwanger and Moretti (2016); Liu et al. (2021), and a light pseudoscalar particle $a$ decaying to $e^{+}e^{-}$ was proposed to explain the ATOMKI anomaly. The ALPs could also cause a deviation from the expected value of the electron anomalous magnetic moment Parker et al. (2018); Morel et al. (2020); Andreev et al. (2021).
The ten billion $J/\psi$ events collected with the BESIII detector during 2009-2019 Ablikim et al. (2022) offer an excellent opportunity to measure the TFF, and to search for $CP$-violation and hadronically coupled ALPs.
@@ II BESIII DETECTOR
The BESIII detector Ablikim et al. (2010) records symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring Yu et al. (2016) in the center-of-mass energy range from 2.0 to 4.95 GeV, with a peak luminosity of $1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.77\;\text{GeV}$. BESIII has collected large data samples in this energy region Ablikim et al. (2020). The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The magnetic field was 0.9 T in 2012, which affects 10% of the total $J/\psi$ data. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1~{}{\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the TOF barrel region is 68 ps, while that in the end cap region was 110 ps. The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps Cao et al. (2020). About 13% of the data used here benefits from this upgrade.
@@ III DATA SAMPLE AND MONTE CARLO SIMULATION
This analysis is based on $(10087\pm 44)\times 10^{6}$ $J/\psi$ events collected with the BESIII detector from 2009 to 2019 Ablikim et al. (2022); it uses the radiative decay $J/\psi\rightarrow\gamma\eta$, resulting in a sample of about $1.1\times 10^{7}$ $\eta$ events.
The estimation of backgrounds and the determination of detection efficiencies are performed with Monte Carlo (MC) simulations. The BESIII detectors are modeled with GEANT4 Agostinelli et al. (2003); Allison et al. (2006); Huang et al. (2022). The production of the $J/\psi$ resonance is implemented with the MC event generator KKMC Jadach et al. (2000, 2001), while the decays are simulated by EVTGEN Ping (2008); Lange (2001). The possible backgrounds are studied using a sample of $J/\psi$ inclusive events in which the known decays of $J/\psi$ are modeled with branching fractions set to the world average values in the PDG Navas et al. (2024), while the unknown decays are generated with the LUNDCHARM model Chen et al. (2000). For this analysis, specific generators for the $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ Zhang et al. (2012), $\eta\rightarrow\pi^{+}\pi^{-}\pi^{0}$ Qin et al. (2018), $\eta\rightarrow\gamma\pi^{+}\pi^{-}$ Qin et al. (2018), $\eta\rightarrow\gamma e^{+}e^{-}$ Qin et al. (2018), $\eta^{\prime}\rightarrow\eta\pi^{+}\pi^{-}$ Qin et al. (2018) decays are developed based on theoretical amplitudes.
@@ IV Event selection and background analysis
For the decay $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$, a requirement of $\chi^{2}_{\rm{sum}}(\pi^{+}\pi^{-}e^{+}e^{-})<50$ is imposed. The primary peaking background comes from the $\eta\rightarrow\gamma\pi^{+}\pi^{-}$ events, where the photon converts to an $e^{+}e^{-}$ pair. To exclude these background events, a photon conversion finder algorithm Xu and He (2012) is applied to the selected $e^{+}e^{-}$ pair. The distribution of $R_{xy}$, defined as the projected distance from the $e^{+}e^{-}$-vertex position to the IP in the $x-y$ plane Xu and He (2012), is shown in Fig. 1.
Photon conversion events are identified based on their proximity to the beam pipe ($R_{xy}\approx 3.5$ cm) and the inner wall of the MDC ($R_{xy}\approx 6.5$ cm). The observable $\Phi_{ee}$ is the lab-frame opening angle of the $e^{+}e^{-}$ pair Adare et al. (2010), and the two-dimensional distribution of $\Phi_{ee}$ vs. $R_{xy}$ is shown in Fig. 2. For photon conversion events, $\Phi_{ee}$ is expected to be close to zero. After the application of the photon conversion veto $\Phi_{ee}<\frac{\pi}{3}$ when $2.0\ \mathrm{cm}<R_{xy}<7.5\ \mathrm{cm}$, most conversion events are rejected. The remaining background events from the inclusive MC sample are studied, and the dominant background channels listed in Table 2. Dedicated exclusive MC samples are then generated to better estimate their contributions.
@@
Table 2: Main background processes and normalized yields of the $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay. For the $\mathcal{B}$ determination, “--” indicates that the number of background events are float in the fit. “Other” refers to the analyses for TFF, $\mathcal{A}_{CP}$, and an ALP. A more detailed explanation is provided in the subsequent sections.
@@
For the decay $\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$, a requirement of $\chi^{2}_{\rm{sum}}(\pi^{+}\pi^{-}\mu^{+}\mu^{-})<50$ is imposed. The signal window is determined by fitting the $\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ invariant mass spectrum of the signal MC sample. The signal shape is represented by a double Gaussian function, and the signal window is defined as $[\mu-3\sigma,\mu+3\sigma]$, where $\mu$ and $\sigma$ are the weighted mean and standard deviation from the fit. In this case, the signal window is set to $0.531\ \mathrm{GeV}/c^{2}<M(\pi^{+}\pi^{-}\mu^{+}\mu^{-})<0.567\ \mathrm{GeV}/c^{2}$, and there are no events in this $\eta$ mass region, as displayed in Fig. 3.
@@ V.1 Branching Fraction Measurement
To extract the number of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ events, an unbinned maximum likelihood fit is performed to the $\pi^{+}\pi^{-}e^{+}e^{-}$ invariant mass spectrum. The signal and all the background shapes are taken from the MC simulations. The number of background events from $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ is determined from the fit, and all other background events are fixed according to the branching fractions from the PDG Navas et al. (2024), as summarized in Table 2. The fit yields $N_{\rm{sig}}=680\pm 27$ signal events, and the result of the invariant mass of $\pi^{+}\pi^{-}e^{+}e^{-}$ is shown in Fig. 4. The fit $\chi^{2}$ per degree of freedom ($ndf$) is $\chi^{2}/ndf=75.1/98$. The branching fraction is determined by
@@
Here $N_{J/\psi}$ is the number of $J/\psi$ events Ablikim et al. (2022) and $\mathcal{B}(J/\psi\rightarrow\gamma\eta)$ is the branching fraction of $J/\psi\rightarrow\gamma\eta$ Navas et al. (2024). With a detection efficiency of $\varepsilon=(20.12\pm 0.04)\%$, the branching fraction of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ is calculated to be $(3.07\pm 0.12)\times 10^{-4}$, where the uncertainty is statistical only.
@@ V.2 Form Factor Measurement
where $i$ is $e\ \mathrm{or}\ \pi$, $\Gamma_{\rho(770)}=149.1$ MeV is the width of $\rho(770)$, and $\Theta$ is the Heaviside step function. By adjusting the values of the $c_{i}$-parameters Benayoun et al. (2010), one can switch between the various VMD models: (I) hidden gauge model ($c_{1}-c_{2}=c_{3}=1$); (II) full VMD model ($c_{1}-c_{2}=1/3,c_{3}=1$);
To extract the parameter $m_{V}$ in the VMD factor, an unbinned maximum likelihood fit to the $e^{+}e^{-}$ and $\pi^{+}\pi^{-}$ invariant mass spectra in data is performed using MINUIT James and Roos (1975). In the fit, $M(\pi^{+}\pi^{-}e^{+}e^{-})$ is required to be in the $\eta$ mass region of $(0.53-0.57)\mathrm{GeV}/c^{2}$. The backgrounds are generated by MC simulation with specific generators as described in Sec. III. The number of background events of $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ is obtained by fitting the $\pi^{+}\pi^{-}e^{+}e^{-}$ invariant mass spectrum, and other background yields, listed in Table 2, are estimated using PDG branching fractions Navas et al. (2024). The probability of observing the $i^{th}$ event characterized by the measured four-momenta $\xi_{i}$ of final-state particles is
@@
$-\omega_{\mathrm{bkg2}}\sum^{N_{\mathrm{bkg2}}}_{k=1}\ln{\mathcal{P}(\xi_{k})}-...\,],$
@@
The fit results of Model I and Model II are both in good agreement with data; Fig. 5 shows the $\pi^{+}\pi^{-}$ and $e^{+}e^{-}$ invariant mass spectrum fits. The results for $m_{V}$ in Model I and Model II are $m_{V}=749\pm 54\ \mathrm{MeV}/c^{2}$ and $m_{V}=748\pm 53\ \mathrm{MeV}/c^{2}$, respectively. The $\chi^{2}/ndf(\pi^{+}\pi^{-},e^{+}e^{-})$ are calculated to be (I) $73.6/41,22.6/36$ and (II) $72.9/41,22.6/36$. A larger $\eta$ data sample, potentially obtainable from experiments like the Super $\tau-$Charm Facility Achasov et al. (2024), is necessary to achieve a more precise measurement.

### v2

@@ Abstract
[…]
@@ I INTRODUCTION
[…]
@@
SM Zillinger et al. (2022) $2.65\pm 0.17$ $6.36\pm 0.39$
CMD-2 Akhmetshin and others (2001) $3.7^{+2.5}_{-1.8}\pm 3.0$ ...
WASA Berlowski and others (2008) $4.3^{+0.2}_{-1.6}\pm 0.4$ $3.6\times 10^{5}$
KLOE Ambrosino and others (2009) $2.68\pm 0.09\pm 0.07$ ...
In the year 2000, the KTeV experiment observed a significant $CP$ asymmetry in the distribution of the T-odd angle $\varphi$ in $K^{0}_{L}\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay Alavi-Harati and others (2000), where $\varphi$ is the angle between the $e^{+}e^{-}$ decay plane and the $\pi^{+}\pi^{-}$ decay plane in the $K^{0}_{L}$ center-of-mass system. However, so far there is no experimental indication of $CP$-violation in flavor-conserving reactions. The $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay provides an opportunity to search for an analogous $CP$-violation asymmetry in the angle distribution between the $e^{+}e^{-}$ and $\pi^{+}\pi^{-}$ decay planes. This has been explored in $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay in the KLOE experiment, and the $CP$-violation asymmetry is determined to be $\mathcal{A}_{CP}=(-0.6\pm 2.5\pm 1.8)\%$ Ambrosino and others (2009), which is consistent with zero.
[…]
The ten billion $J/\psi$ events collected with the BESIII detector during 2009-2019 Ablikim and others (2022) offer an excellent opportunity to measure the TFF, and to search for $CP$-violation and hadronically coupled ALPs.
@@ II BESIII DETECTOR
The BESIII detector Ablikim and others (2010) records symmetric $e^{+}e^{-}$ collisions provided by the BEPCII storage ring Yu and others (2016) in the center-of-mass energy range from 2.0 to 4.95 GeV, with a peak luminosity of $1\times 10^{33}\;\text{cm}^{-2}\text{s}^{-1}$ achieved at $\sqrt{s}=3.77\;\text{GeV}$. BESIII has collected large data samples in this energy region Ablikim and others (2020). The cylindrical core of the BESIII detector covers 93% of the full solid angle and consists of a helium-based multilayer drift chamber (MDC), a plastic scintillator time-of-flight system (TOF), and a CsI(Tl) electromagnetic calorimeter (EMC), which are all enclosed in a superconducting solenoidal magnet providing a 1.0 T magnetic field. The magnetic field was 0.9 T in 2012, which affects 10% of the total $J/\psi$ data. The solenoid is supported by an octagonal flux-return yoke with resistive plate counter muon identification modules interleaved with steel. The charged-particle momentum resolution at $1~{\rm GeV}/c$ is $0.5\%$, and the ${\rm d}E/{\rm d}x$ resolution is $6\%$ for electrons from Bhabha scattering. The EMC measures photon energies with a resolution of $2.5\%$ ($5\%$) at $1$ GeV in the barrel (end cap) region. The time resolution in the TOF barrel region is 68 ps, while that in the end cap region was 110 ps. The end cap TOF system was upgraded in 2015 using multigap resistive plate chamber technology, providing a time resolution of 60 ps Cao and others (2020). About 87% of the data used here benefits from this upgrade.
@@ III DATA SAMPLE AND MONTE CARLO SIMULATION
This analysis is based on $(10087\pm 44)\times 10^{6}$ $J/\psi$ events collected with the BESIII detector from 2009 to 2019 Ablikim and others (2022); it uses the radiative decay $J/\psi\rightarrow\gamma\eta$, resulting in a sample of about $1.1\times 10^{7}$ $\eta$ events.
The estimation of backgrounds and the determination of detection efficiencies are performed with Monte Carlo (MC) simulations. The BESIII detectors are modeled with GEANT4 Agostinelli and others (2003); Allison and others (2006); Huang et al. (2022). The production of the $J/\psi$ resonance is implemented with the MC event generator KKMC Jadach et al. (2000, 2001), while the decays are simulated by EVTGEN Ping (2008); Lange (2001). The possible backgrounds are studied using a sample of $J/\psi$ inclusive events in which the known decays of $J/\psi$ are modeled with branching fractions set to the world average values in the PDG Navas and others (2024), while the unknown decays are generated with the LUNDCHARM model Chen et al. (2000). For this analysis, specific generators for the $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ Zhang et al. (2012), $\eta\rightarrow\pi^{+}\pi^{-}\pi^{0}$ Qin et al. (2018), $\eta\rightarrow\gamma\pi^{+}\pi^{-}$ Qin et al. (2018), $\eta\rightarrow\gamma e^{+}e^{-}$ Qin et al. (2018), $\eta^{\prime}\rightarrow\eta\pi^{+}\pi^{-}$ Qin et al. (2018) decays are developed based on theoretical amplitudes.
@@ IV Event selection and background analysis
For each candidate, particle identification (PID) is performed using the TOF and dEE/dxx information, and a four-constraint (4C) kinematic fit is executed imposing energy and momentum conservation under the hypothesis of $\gamma\pi^{+}\pi^{-}e^{+}e^{-}$ or $\gamma\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ final state. A summed chi-square, $\chi^{2}_{\rm{sum}}=\chi^{2}_{\rm{4C}}+\sum_{i=1}^{4}\chi^{2}_{{\rm PID}(i)}$, is calculated from the chi-square of the 4C kinematic fit ($\chi^{2}_{\rm{4C}}$) and PID ($\chi^{2}_{\rm{PID}}$). For each event, the hypothesis with the smallest $\chi^{2}_{\rm{sum}}(\pi^{+}\pi^{-}e^{+}e^{-})$ or $\chi^{2}_{\rm{sum}}(\pi^{+}\pi^{-}\mu^{+}\mu^{-})$ is kept for the further analysis.
For the decay $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$, a requirement of $\chi^{2}_{\rm{sum}}(\pi^{+}\pi^{-}e^{+}e^{-})<50$ is imposed. The primary peaking background comes from the $\eta\rightarrow\gamma\pi^{+}\pi^{-}$ events, where the photon converts to an $e^{+}e^{-}$ pair. To exclude these background events, a photon conversion finder algorithm Xu and He (2012) is applied to the selected $e^{+}e^{-}$ pair. The distribution of $R_{xy}$, defined as the projected distance from the $e^{+}e^{-}$-vertex position to the IP in the xyx-y plane Xu and He (2012), is shown in Fig. 1.
@@
Table 2: Main background processes and normalized yields of the $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay. For the $\mathcal{B}$ determination, “-” indicates that the number of background events are float in the fit. “Other” refers to the analyses for TFF, $\mathcal{A}_{CP}$, and an ALP. A more detailed explanation is provided in the subsequent sections.
@@
[…]
@@ V.1 Branching Fraction Measurement
To extract the number of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ events, an unbinned maximum likelihood fit is performed to the $\pi^{+}\pi^{-}e^{+}e^{-}$ invariant mass spectrum. The signal and all the background shapes are taken from the MC simulations. The number of background events from $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ is determined from the fit, and all other background events are fixed according to the branching fractions from the PDG Navas and others (2024), as summarized in Table 2. The fit yields $N_{\rm{sig}}=680\pm 27$ signal events, and the result of the invariant mass of $\pi^{+}\pi^{-}e^{+}e^{-}$ is shown in Fig. 4. The fit $\chi^{2}$ per degree of freedom (ndfndf) is $\chi^{2}/ndf=75.1/98$. The branching fraction is determined by
@@
Here $N_{J/\psi}$ is the number of $J/\psi$ events Ablikim and others (2022) and $\mathcal{B}(J/\psi\rightarrow\gamma\eta)$ is the branching fraction of $J/\psi\rightarrow\gamma\eta$ Navas and others (2024). With a detection efficiency of $\varepsilon=(20.12\pm 0.04)\%$, the branching fraction of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ is calculated to be $(3.07\pm 0.12)\times 10^{-4}$, where the uncertainty is statistical only.
@@ V.2 Form Factor Measurement
[…]
To extract the parameter $m_{V}$ in the VMD factor, an unbinned maximum likelihood fit to the $e^{+}e^{-}$ and $\pi^{+}\pi^{-}$ invariant mass spectra in data is performed using MINUIT James and Roos (1975). In the fit, $M(\pi^{+}\pi^{-}e^{+}e^{-})$ is required to be in the $\eta$ mass region of $(0.53-0.57)\mathrm{GeV}/c^{2}$. The backgrounds are generated by MC simulation with specific generators as described in Sec. III. The number of background events of $J/\psi\rightarrow\gamma\pi^{+}\pi^{-}\pi^{+}\pi^{-}$ is obtained by fitting the $\pi^{+}\pi^{-}e^{+}e^{-}$ invariant mass spectrum, and other background yields, listed in Table 2, are estimated using PDG branching fractions Navas and others (2024). The probability of observing the $i^{th}$ event characterized by the measured four-momenta $\xi_{i}$ of final-state particles is
@@
$-\omega_{\mathrm{bkg2}}\sum^{N_{\mathrm{bkg2}}}_{k=1}\ln{\mathcal{P}(\xi_{k})}-$.\,],
@@
The fit results of Model I and Model II are both in good agreement with data; Fig. 5 shows the $\pi^{+}\pi^{-}$ and $e^{+}e^{-}$ invariant mass spectrum fits. The results for $m_{V}$ in Model I and Model II are $m_{V}=749\pm 54\ \mathrm{MeV}/c^{2}$ and $m_{V}=748\pm 53\ \mathrm{MeV}/c^{2}$, respectively. The $\chi^{2}/ndf(\pi^{+}\pi^{-},e^{+}e^{-})$ are calculated to be (I) $73.6/41,22.6/36$ and (II) $72.9/41,22.6/36$. A larger $\eta$ data sample, potentially obtainable from experiments like the Super $\tau$-Charm Facility Achasov and others (2024), is necessary to achieve a more precise measurement.